An asset has four possible rates of return, all equally likely: 5.6%, 15.3%, 15.8%, and 19.8%. What is the standard deviation of this asset's rate of return?
Answer #1
Return(%) Deviation form average return (D) D^2
5.6 -8.53 72.68
15.3 1.18 1.38
15.8 1.68 2.81
19.8 5.68 32.21

Average Return = Sum of Returns / No. of Returns

= (5.6+15.3+15.8+19.8)/4

= 56.5/4

= 14.125%

Variance = \sumD^2/n

= (72.68+1.38+2.81+32.21)/4

= 109.07/4

= 27.27

Standard Deviation = \sqrt{}Variance

= \sqrt{}27.27

= 5.22%

*Deviation form average return = Rate of return -  average return